Suppose the lengths of human pregnancies are normally distributed with 266 days and ơ16 days. Complete parts a and b below. The f ure to he ri ht represents the normal curve with μ-2 6 days and : 16 days. The area to the right O X-290 is 0.0668, provide two interpretations of this area. a Provide one interpretation of the area. Select the correct choice below and fill in the answer boxes to complete your choice. Type integers or decimals. Do not round.) () A. The proportion of human pregnancies that last more than□ days is OB. The proportion of human pregnances that last less than daysǐ
Answer #1

=266 and σ= 16 The area to the right ofx=290 is 0.0668 A) The proportion ofhuman pregnancies last more than 290 days is x-μ P (x >290)- PH 290-u, - 290-266 P(Z> =P(Z > 1.5) 16 0.5-0.4332 - 0.0668 (From Normal table Hence, the proportion of human pregnancies last more than 290 days is 0.0668 B) The proportion ofhuman pregnancies last less than 290 days is P(Z< 2901666) = P(Z < 1.5) P-00Z0) +P0KZ <1.5) = 0.5 0.4332 (From Normal table = 0.9332 Hence, the proportion of human pregnancies last less than 290 days is 0.9332